OpenStudy (anonymous):

An airplane flew 340 miles in 2 hours with the wind. It took 2 and a half hours to make the return trip against the wind. How fast is the wind blowing? I have the formula d=rt and I started a table, but not exactly sure what equation type would give me the right answer.

OpenStudy (anonymous):

d = rt 340 = (r + w)2 340 = (r-w)2.5 distribute both equations 340 = 2r + 2w 340 = 2.5r - 2.5w Solve the 1st one for r and then substitute 340 - 2w = 2r divide by 2 170 - w = r substitute 340 = 2.5(170 - w) - 2.5w 340 = 425 - 2.5w - 2.5w 340 = 425 - 5w -85 = -5w 17 = w wind is 17 mph

OpenStudy (anonymous):

Thnk you very much :D